​(Discounted payback period​) ​ Gio's Restaurants is considering a project with the following expected cash​ flows:

Year

Project Cash Flow​ (millions)

0

​$(180​)

1

100

2

65

3

100

4

110

If the​ project's appropriate discount rate is 13 ​percent, what is the​ project's discounted payback​ period?

The​ project's discounted payback period is _____ years.  ​(Round to two decimal​ places.)
Answer #1
Please find the answer below
Statement showing computation:
Calculation of discounted payback period:
Discounted PBP
Time Amount PVf PV Cumulative
                                                                                      -                          (180.00)                                                    1.0000                                         (180.00)                 (180.00)
                                                                                 1.00                          100.00                                                    0.8850                                              88.50                    (91.50)
                                                                                 2.00                             65.00                                                    0.7831                                              50.90                    (40.60)
                                                                                 3.00                          100.00                                                    0.6931                                              69.31                      28.71
                                                                                 4.00                          110.00                                                    0.6133                                              67.47                      96.17
PBP = 2+40.60/69.31
PBP = 2+0.59= 2.59
Discounted payback period is 2.59 years